lim ( tan x) = ∞

x→π/2

True or false with full explanation


1
Expert's answer
2021-06-28T03:50:31-0400


Taking into account that "\lim\limits_{x\to\frac{\pi}{2}}\sin x=\sin\frac{\pi}{2}=1" and "\lim\limits_{x\to\frac{\pi}{2}}\cos x=\cos\frac{\pi}{2}=0," we conclude that "\lim\limits_{x\to\frac{\pi}{2}}\tan x=\lim\limits_{x\to\frac{\pi}{2}}\frac{\sin x}{\cos x}=|\frac{1}{0}|=\infty."


Answer: true